OpenStudy (anonymous):

When the equation 27 to the power of x end power equals root of order three of 27 squared is evaluated for x, you derive an equivalent form of the radical root of order three of 27 squared in rational exponent form: 27x. What is the value of x? x = 3 x = 2 x = 3 over 2 x = 2 over 3

OpenStudy (anonymous):

which answer is it

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

this is why mathematicians invented notation this is pretty much impossible to decipher try the equation tool below o

OpenStudy (anonymous):

so you dont know the answer

OpenStudy (misty1212):

actually i don't know the question

OpenStudy (anonymous):

what

OpenStudy (misty1212):

is it something like \[27^x\]?

OpenStudy (misty1212):

or maybe \[\huge 27^x=\sqrt[3]{27^2}\]

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

ok now i understand

OpenStudy (misty1212):

\[\huge \sqrt[n]{b^m}=b^\frac{m}{n}\]

OpenStudy (misty1212):

so \[\huge\sqrt[3]{27^2}=27^{\frac{2}{3}}\]

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yeah Probably 2/3

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (misty1212):

it is \[\frac{2}{3}\] yes

OpenStudy (misty1212):

the actual number, however, is \(3^2=9\)